# Advancing Glitch Classification in Gravity Spy: Multi-view Fusion with Attention-based Machine Learning for Advanced LIGO’s Fourth Observing Run

### v1

@@ 3.4 Effects of Gated Attention Mechanism
Table 3 compares the performance of models with and without the gated attention mechanism. These models share an identical CNN architecture for feature extraction from each view but diverge in their feature fusion approaches, such as mean pooling and max pooling. The results demonstrate that the performance of the model with the gated attention mechanism is significantly better than of those without it. Additionally, statistical analysis indicates no significant difference between the performances of mean pooling and max pooling.
The gated attention mechanism also enhances interpretability, rendering the model’s outputs more explainable to humans. Figure 6 shows examples with attention weights assigned by the proposed classifier for four glitch predictions. The magnitude of attention weights positively correlates with the contribution of each time view to the prediction. For instance, in Fig. 6 (a), (d) and (e), the model places greater emphasis on long time windows to predict glitches like Repeating Blips and No Glitch, while allocating higher attention to short time windows to predict glitches such as Blip and Koi Fish in Fig. 6 (b) and (c).
Table 3: Model evaluations (the overall mean ± standard deviation across 5 independent runs) for classifiers with gated attention mechanism and without attention mechanism (using max pooling and mean pooling) on the test set. Best results are highlighted in bol, and * indicates that the comparison is statistically significant ($p<0.05$).
Attention
Mechanism
Proposed
(Gated Attention)
No Attention
(Max Pooling )
No Attention
(Mean Pooling)
@@
Figure 6: Illustrative examples of glitch classification with attention weights generated by the classifier for each time window view. Attention weights are shown in white in each subplot. A higher attention weight indicates that the model relies more on that particular view for predicting the glitch class.
@@ 3.5 Comparison with Previous Classifiers
We compare the performance of the proposed model with the previously implemented classifier soni2021discovering in the O3 run on the same test set, as shown in Table 4. The proposed model shows superiority across all metrics, outperforming the previous model by 2.7% to 3.7%. Furthermore, a detailed examination of class-wise accuracies reveals that the proposed model provides more accurate predictions for the majority of glitch classes. In particular, the model excels in classes like Wandering Line, Repeating Blips, and Helix. While there are a few instances for which the previous classifier attains better outcomes, such as Koi Fish and Scattered Light, the differences remain relatively modest.
@@
Figure 7: Comparison of t-SNE plots in a two-dimensional space (latent feature $z_{1}$ and $z_{2}$) between the previous classifier and the newly proposed classifier on the same test set. The proposed classifier (right) generates more tightly-clustered glitch features than the previous one (left) within the latent space.
@@
To obtain a deeper insight into this finding, we compare the t-SNE plots of these two classifiers within a two-dimensional space, as shown in Fig. 7. Compared to the previous classifier, it is evident that the feature distributions generated by the new classifier exhibit tighter clusters within the same glitch class and distinct separation among clusters of different classes. However, some degree of overlap persists between glitch classes with similar morphologies, such as Extremely Loud and Koi Fish, Blip and Repeating Blips, Scattered Light and Low Frequency Lines.
@@
Figure 8: The reliability plots of the model output probabilities on the test set. The plots illustrate the observed fraction of positives against the predicted fraction of positives. The diagonal dotted line indicates perfect reliability. The expected calibration error (ECE) quantifies the average discrepancy between predicted probabilities and actual accuracy.
@@
Figure 8 presents the reliability plots of two classifiers applied to the test set. Both classifiers are overconfident in their predictions, evidenced by the majority of points residing below the diagonal line. Nevertheless, the proposed model demonstrates better calibration compared to the previous model, particularly in the high confidence range ($>0.8$) – a crucial aspect for detector characterization and gravitational-wave data analysis. Additionally, the proposed model achieves a lower expected calibration error of 0.06, compared to the previous model’s 0.09. This discrepancy indicates that the proposed model displays less bias and higher reliability than the previous classifier.
@@
Table 4: Comparison of class-wise classification accuracy and the overall performance (accuracy, precision, F1 and AUC scores) between the previous classifier and the proposed classifier. Better results are highlighted in bold, and * indicates that the comparison is statistically significant ($p<0.05$).
Class Previous Proposed Class Previous Proposed 1080Lines 0.985 0.991 Paired Doves 0.849 0.882* 1400Ripples 0.967 0.979* Power Line 0.987 0.994* Air Compressor 0.921 0.954* Repeating Blips 0.772 0.933* Blip 0.980 0.991* Scattered Light 0.929 0.924 Blip Low Frequency 0.926 0.928 Scratchy 0.862 0.957* Chirp 0.913 0.926* Tomte 0.963 0.984* Extremely Loud 0.901 0.960* Violin Mode 0.924 0.960* Fast Scattering 0.884 0.890 Wandering Line 0.52 0.740* Helix 0.732 0.891* Whistle 0.957 0.972* Koi Fish 0.921 0.917 Overall Light Modulation 0.918 0.947* Accuracy 0.906 0.941* Low Frequency Burst 0.962 0.981* Precision 0.913 0.950* Low Frequency Lines 0.818 0.916* F1 0.904 0.931* No Glitch 0.978 0.984 AUC 0.937 0.965*
@@ 4 Discussion
In this study, we introduce a novel classifier for the Gravity Spy project, aiming to enhance the accuracy of glitch classification provided by the LIGO gravitational-wave detectors. Overall, our proposed classifier achieves accuracy of 0.941, precision of 0.950, F1 score of 0.931 and AUC of 0.965 on the held-out set of glitches, outperforming prior classifiers integrated into the previous Gravity Spy workflow.
The proposed classifier exhibits superior glitch classification through various attributes. First, the use of the inception residual block empowers the model to deepen its architecture without overfitting concerns, thereby facilitating the extraction of more discriminative features unique to different glitch classes. The tightly-clustered distributions in Fig. 7 further substantiate this finding. Next, the intermediate fusion strategy outperforms early and late fusion due to its ability to capture and combine cross-view correlations, which aligns with findings from prior studies, such as using multiple time scales for anomaly detection wang2021intermediate and multiple modalities for gesture recognition roitberg2019analysis . Moreover, label smoothing effectively tackles the issue of noisy labels within the training dataset by introducing a controlled level of uncertainty into label distributions during training, resulting in enhanced glitch predictions. However, it is crucial to choose the level of this uncertainty infusion appropriately, as excessive smoothing, illustrated in Fig. 5, can lead to label confusion within the model, hindering its ability to learn effectively.
@@
In our study, the model achieves interpretability by discerning which view contributes the most to the classification decision. Consequently, the attention mechanism significantly improves the overall model performance by selectively prioritizing informative aspects of each view and effectively capturing cross-view relationships. For example, as shown in Fig. 6, the model’s extraction of more discriminative features from the 4 s and 2 s windows when classifying a glitch as Repeating Blips aligns with volunteers’ experiences, reflecting their awareness that repeating patterns primarily manifest in longer time windows. In contrast, when identifying a single Blip class, the model focuses on extracting more specific features from the 0.5 s window, corresponding to the transient nature of this particular glitch category. Similarly, when dealing with cases resembling Blip or Koi Fish, the model directs greater attention to the 0.5 s window. Concerning No Glitch classification, the model’s emphasis on the 4 s window is intuitive, aligning with the necessity to inspect the longest window time to confirm all background features or, in other words, the absence of glitches. It is essential to note that the emphasis on specific windows does not diminish the importance of those with low attention weights. Instead, the model effectively captures crucial information from all four windows, enabling it to make optimal glitch predictions. Potentially, the attention mechanism can guide volunteers and LIGO experts, enabling them to enhance their understanding of glitches and thereby facilitating a more effective characterization process.
@@
Figure 9: Examples of glitches with different predicted classes between the previous model and the proposed model. (a) Initially predicted as the Koi Fish class by the previous model, it is now classified as Extremely Loud by the proposed model. (b) Initially classified as the Scattered Light class by the previous model, it is now predicted as Low Frequency Lines by the proposed model.
@@
Given its superior overall performance in comparison to the previous classifier, the proposed classifier has been integrated into the Gravity Spy system for the O4 run. This integration coincides with major infrastructure enhancements introduced in the O4 run cahillane2022review ; abbott2020prospects , aimed at achieving higher instrument sensitivity for the detection of gravitational-wave signals. However, this improvement in detector sensitivity is accompanied by a trade-off, i.e., a higher sensitivity to environmental and instrumental artifacts, which potentially leads to new glitch classes. In this context, the more discriminative features extracted by the proposed classifier, as shown in Fig. 7, hold the potential to more effectively identify and categorize these new glitch classes.
@@
An underperformance is observed in certain glitch classes, such as Koi Fish and Scattered Light as presented in Table 4. It is essential to note that this underperformance is not significant ($p>0.05$), approximately $0.5\%$ compared to the previous model, which may potentially be attributed to random sampling variations in our test set. Besides, this underperformance is also influenced by the inherent complexity and variability within these classes. Given their prevalence and diverse morphologies, establishing strict boundaries for these classes poses a challenge, leading to expected variations in classifiers when distinguishing these intricacies. For instance, in Fig. 9, Koi Fish can be confused with Extremely Loud when the trigger is very loud soni2021discovering and Scattered Light can be confused with Low Frequency Lines as they share characteristics of low-frequency and long duration accadia2010noise .
@@
Figure 10: The density distribution plot of predicted confidence for O4 glitches between (a) the previous model and (b) the proposed model.
@@
To demonstrate the impact of confidence overfitting in the O4 run, we conduct a comparative analysis of the predicted confidence distributions between the previous model and the proposed model. This analysis encompasses a total of 2267 glitches extracted from May 24, 2023, to September 19, 2023, within the O4 run. As illustrated in Fig. 10 (a), the previous model exhibits persistent confidence overfitting, as most glitch classes receive predictions with confidence close to 1, except for Paired Doves and Light Modulation. In contrast, the proposed model, as shown in Fig. 10 (b), significantly mitigates this issue, providing predictions with varying confidence ranges across glitch classes. Specifically, for glitch classes like Extremely Loud and Low Frequency Lines, which exhibit less variation in morphologies within the O4 distribution compared to O3, the proposed model confidently expresses predictions with relatively high confidence. Conversely, for glitch classes such as Whistle, Wandering Line, Violin Mode, and Scratchy, the proposed model displays lower confidence in predictions, given the visually distinct morphologies in O4 compared to their distributions in O3. This distinction is crucial for LIGO experts during glitch analysis and characterization, allowing the exclusion of glitches with lower confidences to enhance the precision of their analyses. This, in turn, opens avenues to explore the potential discovery of new glitch classes within clusters exhibiting lower confidences. A good example is the identification of a new glitch class, named 589 Hz,  [^2] https://www.zooniverse.org/projects/zooniverse/gravity-spy/talk/762/2986261 depicted in Fig. 11, which occurred on May 30, 31, and June 1 of 2023 in the Hanford detector. The previous model predicted it as Whistle with an overall high confidence of 0.976(±0.0590.976($\pm$ 0.0590.976 ( ± 0.059) on 267 collections, attributing this classification to its high-frequency nature similar to Whistle, while the proposed model predicted it with significantly lower confidence scores of 0.463(±0.1380.463($\pm$ 0.1380.463 ( ± 0.138) for 195 cases classified as Whistle, $0.428(\pm 0.137)$ for 49 cases classified as No Glitch and $0.347(\pm 0.057)$ for 23 cases classified as Violin Mode, respectively.
@@
Figure 11: An example of a new glitch class, named 589 Hz in the O4 run. It is a narrow band line glitch near 589 Hz lasting less than 0.4 s, appearing in these spectrograms as bright yellow for 0.02 s to 0.2 s. It was classified as the Whistle class with a high confidence of 0.972 by the previous classifier and a low confidence of 0.348 by the proposed classifier.
@@
This study has some limitations. First, as the classifier is trained under full supervision, its generalization capability to new classes is constrained. Consequently, if a novel glitch class arises during the O4 run, the classifier requires retraining on a dataset that incorporates the new glitch class. This process also involves the participation of volunteers who, with assistance from the Gravity Spy similarity search coughlin2019classifying , identify and categorize these new glitch classes. Next, the training and test set used in this study are still imperfect with noisy labels. To improve the reliability of the model, future efforts should be put into preparing a test set with validated ground truth, as confirmed by LIGO experts. Additionally, there is room for further optimization of the classifier for improved glitch classification. Future studies can enhance the model’s training by incorporating additional information such as sensor locations, temporal attributes, and supplementary auxiliary data zevin2023gravity ; nguyen2021environmental ; colgan2020efficient . For instance, supplementary auxiliary channels, could assist in reducing potential Scattered Light, as illustrated in Fig. 9 (b), by aiding in the identification of ground motion soni2020reducing . Last, our model currently provides information on important views but does not specify which parts of the focused image trigger predictions. Our research will explore techniques like Grad-CAM 8237336 to precisely identify the influential regions within glitch images.

### v2

@@ 3.4 Effects of gated attention mechanism
Table 3 compares the performance of models with and without the gated attention mechanism. These models share an identical convolutional neural network architecture for feature extraction from each view but diverge in their feature fusion approaches, such as mean pooling and max pooling. The results demonstrate that the performance of the model with the gated attention mechanism is significantly better than of those without it. Additionally, statistical analysis indicates no significant difference between the performances of mean pooling and max pooling.
The gated attention mechanism also enhances interpretability, rendering the model’s outputs more explainable to humans. Figure 6 shows examples with attention weights assigned by the proposed classifier for four glitch predictions. The magnitude of attention weights positively correlates with the contribution of each time view to the prediction. For instance, in Figure 6 (a), (d) and (e), the model places greater emphasis on long time windows to predict glitches like Repeating Blips and No Glitch, while allocating higher attention to short time windows to predict glitches such as Blip and Koi Fish in Figure 6 (b) and (c).
Table 3: Model evaluations (the overall mean ± standard deviation across five independent runs) for classifiers with gated attention mechanism and without attention mechanism (using max pooling and mean pooling) on the O3 test set using the intermediate fusion structure. Best results are highlighted in bold, and an asterisk (*) indicates that the comparison is statistically significant given our chosen threshold ($p<0.05$).
Attention mechanism
Proposed No Attention No Attention
Performance metric (gated attention) (max pooling) (mean pooling)
@@
Figure 6: Illustrative examples of glitch classification with attention weights generated by the classifier for each time-window view. Attention weights are shown in white in each subplot. A higher attention weight indicates that the model relies more on that particular view for predicting the glitch class.
@@ 3.5 Comparison with previous classifiers
We compare the performance of the proposed model with the previously implemented classifier [24]. The previous classifier has been retrained using the same dataset as our new model to enable a clean comparison. Results on our O3 test set are shown in Table 4. The proposed model shows superiority across all metrics, outperforming the previous model by $2.7\%$ to $3.7\%$. Furthermore, a detailed examination of class-wise accuracies reveals that the proposed model provides more accurate predictions for the majority of glitch classes. In particular, the model performs better for classes like Wandering Line, Repeating Blips, and Helix. While there are a few instances for which the previous classifier attains better outcomes, such as Koi Fish and Scattered Light (Slow Scattering), the differences remain relatively modest.
@@
Figure 7: Comparison of $t$-SNE plots in a two-dimensional space (latent features $t$-SNE 11 and $t$-SNE 22) between the previous classifier and the newly proposed classifier on the same O3 test set). The proposed classifier (right) generates more tightly-clustered glitch features than the previous one (left) within the latent space. Both plots were generated using the same $t$-SNE parameters for consistency.
@@
To obtain a deeper insight into this finding, we compare the $t$-SNE plots of these two classifiers within a two-dimensional space, as shown in Figure 7. Compared to the previous classifier, it is evident that the feature distributions generated by the new classifier exhibit tighter clusters within the same glitch class and distinct separation among clusters of different classes. However, some degree of overlap persists between glitch classes with similar morphologies, such as Extremely Loud and Koi Fish, Blip and Repeating Blips, Scattered Light and Low-frequency Lines.
@@
Figure 8: The reliability plots of the model output probabilities on the O3 test set. The plots illustrate the observed fraction of positives against the predicted fraction of positives. The diagonal dotted line indicates perfect reliability. The expected calibration error (ECE) quantifies the average discrepancy between predicted probabilities and actual accuracy.
@@
Figure 8 presents the reliability plots of two classifiers applied to the test set. Both classifiers are overconfident in their predictions, evidenced by the majority of points residing below the diagonal line. Nevertheless, the proposed model demonstrates better calibration compared to the previous model, particularly in the high confidence range (>0.8>0.8)—a crucial aspect for detector characterization and gravitational-wave data analysis. Additionally, the proposed model achieves a lower expected calibration error of $0.06$, compared to the previous model’s $0.09$. This discrepancy indicates that the proposed model displays less bias and higher reliability than the previous classifier.
@@
Table 4: Comparison of class-wise classification accuracy, and the overall performance (accuracy, precision, F1 and AUC scores) between the previous classifier and the proposed classifier evaluated with our O3 dataset. Better results are highlighted in bold, and an asterisk (*) indicates that the comparison is statistically significant given our chosen threshold ($p<0.05$).
Classifier
Performance metric Previous Proposed
Class accuracy 1080 Lines 0.985 0.991
1400 Ripples 0.967 0.979*
Air Compressor 0.921 0.954*
Blip 0.980 0.991*
Blip Low Frequency 0.926 0.928
Chirp 0.913 0.926*
Extremely Loud 0.901 0.960*
Fast Scattering 0.884 0.890
Helix 0.732 0.891*
Koi Fish 0.921 0.917
Light Modulation 0.918 0.947*
Low-frequency Burst 0.962 0.981*
Low-frequency Lines 0.818 0.916*
No Glitch 0.978 0.984
Paired Doves 0.849 0.882*
Power Line 0.987 0.994*
Repeating Blips 0.772 0.933*
Scattered Light 0.929 0.924
Scratchy 0.862 0.957*
Tomte 0.963 0.984*
Violin Mode 0.924 0.960*
Wandering Line 0.520 0.740*
Whistle 0.957 0.972*
Overall Accuracy 0.906 0.941*
Precision 0.913 0.950*
F1 0.904 0.931*
AUC 0.937 0.965*
@@ 4 Discussion
In this study, we introduce a novel classifier for the Gravity Spy project, aiming to enhance the accuracy of glitch classification provided for LIGO. Overall, our proposed classifier achieves accuracy of $0.941$, precision of $0.950$, F1 score of $0.931$ and AUC of $0.965$ on the held-out set of glitches, outperforming prior classifiers integrated into the previous Gravity Spy workflow.
The proposed classifier exhibits superior glitch classification through various attributes. First, the use of the inception residual block empowers the model to deepen its architecture without overfitting concerns, thereby facilitating the extraction of more discriminative features unique to different glitch classes. The tightly clustered distributions in Figure 7 further substantiate this finding. Next, the intermediate fusion strategy outperforms early and late fusion due to its ability to capture and combine cross-view correlations, which aligns with findings from prior studies, such as using multiple time scales for anomaly detection [57] and multiple modalities for gesture recognition [58]. Moreover, label smoothing effectively tackles the issue of noisy labels within the training dataset by introducing a controlled level of uncertainty into label distributions during training, resulting in enhanced glitch predictions. However, it is crucial to choose the level of this uncertainty infusion appropriately, as excessive smoothing, illustrated in Figure 5, can lead to label confusion within the model, hindering its ability to learn effectively.
@@
In our study, the model achieves interpretability by discerning which view contributes the most to the classification decision. Consequently, the attention mechanism significantly improves the overall model performance by selectively prioritizing informative aspects of each view and effectively capturing cross-view relationships. For example, as shown in Figure 6, the model’s extraction of more discriminative features from the $2~\mathrm{s}$ and $4~\mathrm{s}$ windows when classifying a glitch as Repeating Blips aligns with volunteers’ experiences, reflecting their awareness that repeating patterns primarily manifest in longer time windows. In contrast, when identifying a single Blip class, the model focuses on extracting more specific features from the $0.5~\mathrm{s}$ window, corresponding to the transient nature of this particular glitch category. Similarly, when dealing with cases resembling Blip or Koi Fish, the model directs greater attention to the $0.5~\mathrm{s}$ window. Concerning No Glitch classification, the model’s emphasis on the $4~\mathrm{s}$ window is intuitive, aligning with the necessity to inspect the longest window time to confirm all background features and the absence of glitches. For Whistle, some attention is given to the $0.5~\mathrm{s}$ window, where the detail of the characteristic up-and-down sweep is most visible, but more is given to the $2~\mathrm{s}$ window. While the illustrative Whistle is a single V-shaped sweep, Whistles can often contain multiple sweeps (a W shape as shown in Figure 1 is common), and these are more apparent in the longer windows: the $2~\mathrm{s}$ window seems to give the best balance between resolving the up-and-down structure and the potential repetition of sweeps. The emphasis on specific windows does not diminish the importance of those with low attention weights. Instead, the model effectively captures crucial information from all four windows, enabling it to make optimal glitch predictions. Potentially, the attention mechanism can guide volunteers and LIGO experts, enabling them to enhance their understanding of glitches and thereby facilitating a more effective characterization process.
@@
Figure 9: Examples of glitches with different predicted classes between the previous model and the proposed model. (a) Initially predicted as the Koi Fish class by the previous model, it is now classified as Extremely Loud by the proposed model. (b) Initially classified as the Scattered Light class by the previous model, it is now predicted as Low-frequency Lines by the proposed model.
@@
Given its superior overall performance in comparison to the previous classifier, the proposed classifier has been integrated into the Gravity Spy system for the O4 run. This integration coincides with major infrastructure enhancements introduced in the O4 run [59, 26, 27, 28], aimed at achieving higher instrument sensitivity for the detection of gravitational-wave signals. However, this improvement in detector sensitivity is accompanied by a trade-off, i.e., a higher sensitivity to environmental and instrumental artifacts, which potentially leads to new glitch classes.
In this context, the more discriminative features extracted by the proposed classifier, as shown in Figure 7, have the potential to more effectively identify and categorize these new glitch classes. Additionally, Figure 7 also shows distinct clusters within a single class, which suggests the possible existence of sub-classes that may be interesting to explore
@@
An underperformance is observed in certain glitch classes, such as Koi Fish and Scattered Light as presented in Table 4. However, this underperformance is not significant ($p>0.05$), approximately $0.5\%$ compared to the previous model, which may potentially be attributed to random sampling variations in our test set. Besides, this underperformance is also influenced by the inherent complexity and variability within these classes. Given their prevalence and diverse morphologies, establishing strict boundaries for these classes poses a challenge, leading to expected variations in classifiers when distinguishing these intricacies. For instance, in Figure 9, Koi Fish can be confused with Extremely Loud when the trigger is very loud [24], and Scattered Light can be confused with Low-frequency Lines as they share characteristics of being low frequency and long duration [60].
@@
Figure 10: The density distribution plot of predicted confidence for early O4 glitches between (a) the previous model and (b) the proposed model. The wider sections indicating higher densities. The black horizontal bars within each section represent the median confidence score for that glitch class.
@@
To demonstrate the impact of confidence overfitting in O4, we conduct a comparative analysis of the predicted confidence distributions between the previous model and the proposed model. This analysis encompasses a total of $2267$ glitches extracted from early in O4, from 24 May 2023 to 19 September 19 2023 in Hanford and Livingston data. As illustrated in Figure 10 (a), the previous model exhibits persistent confidence overfitting, as most glitch classes receive predictions with confidence close to $1$, except for Paired Doves and Light Modulation. In contrast, the proposed model, as shown in Figure 10 (b), significantly mitigates this issue, providing predictions with varying confidence ranges across glitch classes. Specifically, for glitch classes like Extremely Loud and Low-frequency Lines, which exhibit less variation in morphologies within the O4 distribution compared to O3, the proposed model confidently expresses predictions with relatively high confidence. Conversely, for glitch classes such as Whistle, Wandering Line, Violin Mode, and Scratchy, the proposed model displays lower confidence in predictions, given the visually distinct morphologies in O4 compared to their distributions in O3. This distinction is crucial for LIGO experts during glitch analysis and characterization, allowing the exclusion of glitches with lower confidences to enhance the precision of their analyses. This, in turn, opens avenues to explore the potential discovery of new glitch classes within clusters exhibiting lower confidences. A good example is the identification of a new glitch class, named 589 Hz,222www.zooniverse.org/projects/zooniverse/gravity-spy/talk/762/2986261 depicted in Figure 11, which occurred on May 30, 31, and June 1 of 2023 in the Hanford detector. The previous model predicted it as Whistle with an overall high confidence of $0.976(\pm 0.059$) on $267$ collections, attributing this classification to its high-frequency nature similar to Whistle, while the proposed model predicted it with significantly lower confidence scores of $0.463(\pm 0.138$) for $195$ cases classified as Whistle, $0.428(\pm 0.137)$ for $49$ cases classified as No Glitch and $0.347(\pm 0.057)$ for $23$ cases classified as Violin Mode, respectively.
@@
Figure 11: An example of a new O4 glitch class, named 589 Hz. This class appears as a narrow-band line glitch near $589~\mathrm{Hz}$ lasting less than $0.4~\mathrm{s}$, appearing in these spectrograms as bright yellow for $0.02~\mathrm{s}$ to $0.2~\mathrm{s}$. This example was classified as the Whistle class with a high confidence of $0.972$ by the previous classifier and a low confidence of $0.348$ by the proposed classifier.
@@
This study has some limitations. First, as the classifier is trained under full supervision, its generalization capability to new classes is constrained. Consequently, if a novel glitch class arises during the O4 run, the classifier requires retraining on a dataset that incorporates the new glitch class. This process also involves the participation of volunteers who, with assistance from tools such as the the Gravity Spy similarity search [19], identify and categorize new glitch classes [24]. Next, the training and test set used in this study are still imperfect with noisy labels. To improve the reliability of the model, future efforts should be put into preparing a test set with validated ground truth, as confirmed by LIGO detector-characterization experts. Additionally, there is room for further optimization of the classifier for improved glitch classification. Future studies can enhance the model’s training by incorporating additional information such as sensor locations, temporal attributes, and supplementary auxiliary data [18, 61, 62, 21]. For instance, supplementary auxiliary channels, could assist in reducing potential Scattered Light, as illustrated in Figure 9 (b), by aiding in the identification of ground motion [63]. Last, our model currently provides information on important views but does not specify which parts of the focused image trigger predictions. Our research will explore techniques like Grad-CAM [64] to precisely identify the influential regions within glitch images.